# Orthogonal frequency-division multiplexing for simultaneous gate operations on multiple qubits via a shared control line

### v1

@@ I Introduction
Our findings expand the design space for spectrum allocation and pulse length of microwave signals in the FDM scheme, facilitating the development of a scalable FDM-based microwave controller for quantum processors with a large number of qubits. Notably, we provide design guidelines for the FDM scheme to suppress interference and achieve high fidelity without requiring additional circuit elements. Moreover, the developed framework enables the use of closely spaced drive frequencies and short pulse lengths, thereby achieving high throughput in gate operations. Our approach is inherently scalable and well suited for controlling a large number of qubits owing to its simplicity and narrow bandwidth requirements.
@@ II Model
Figure 1: Conceptual illustration of a model comprising $N=5$ independent qubits with frequencies $\omega_{\mathrm{q},k}$ driven by microwaves via a shared control line. The controller produces NN microwave tones at distinct frequencies $\omega_{\mathrm{d},k}$, which are then combined and routed to a quantum processor via the shared line. The combined signal is applied uniformly to all qubits.
@@
and $\tau$ is the pulse length. In this context, the qubit index set $K$ is given by $K=\left\{k\in\mathbb{Z}\,|\,l\leq k\leq r\right\}$, where rl+1=Nr-l+1=N.
@@ III General analysis
by using $\mathrm{sinc}\left(x\right)\equiv\sin\left(x\right)/x$. Here, $\gamma\subset K$ is the set of qubit indices that lack corresponding elements related to $k_{0}$ (see Fig. 2).
@@
$\lambda_{\mathrm{z}}$ $\equiv\frac{\alpha^{2}\tau}{4}\left[\sum_{k\in\gamma}\frac{\cos(\Delta_{kk_{0}}\tau)-\frac{3}{2}\mathrm{sinc}(\Delta_{kk_{0}}\tau)}{\Delta_{kk_{0}}}\right.+\sum_{k\in\gamma}\sum_{\begin{subarray}{c}j\in K\\ j\neq k_{0}\end{subarray}}\left\{\frac{\mathrm{sinc}\left(\Delta_{jk_{0}}\tau\right)}{\Delta_{kk_{0}}}+\frac{\mathrm{sinc}\left[\left(\Delta_{kk_{0}}+\Delta_{jk_{0}}\right)\tau\right]}{2\Delta_{jk_{0}}}\right\}$
@@ IV Case studies
Figure 3: Absolute values of the spectra of $s\left(t\right)\sin\left(\omega_{\mathrm{d},k}t\right)$ for $N=5$ and $k\in\left\{-2,-1,0,1,2\right\}$, with pulse length (a) $\tau=\tau_{0}$, (b) $\tau=\tau_{0}/2$, and (c) $\tau=2\tau_{0}$. In (a) and (c), each drive frequency $\omega_{\mathrm{d},k}$ is placed at the zero-crossing point of the spectral components of all other microwaves.
Refer to caption Refer to caption Refer to caption
@@ IV.3 Long-pulse limit ($\tau\gg\tau_{0}$)
IV.4 Discussion
The pulse length is a crucial factor for enabling dense spacing of drive frequencies while maintaining high gate fidelity. From the results presented in Secs. IV.1 and IV.2, we observe that high fidelity can be attained near $k_{0}=0$ at both $\tau=\tau_{0}$ and $\tau=\tau_{0}/2$. This is because the rotation angles are $\lambda_{\mathrm{x}}\sim\phi/2$ and $\lambda_{\mathrm{y}},\lambda_{\mathrm{z}}\sim 0$ in each case, which is necessary to achieve high fidelity. An equivalent result can be obtained for every integer multiple of $\tau_{0}$ and $\tau_{0}/2$. Consequently, we conclude that $\tau$ should be chosen from the set $\left\{m\tau_{0}/2\right\}$, where $m$ is a positive integer, to obtain higher fidelity around $k_{0}=0$. In addition, a larger $\tau$ guarantees high fidelity because the rotation angles approach $\lambda_{\mathrm{x}}\rightarrow\phi/2$ and $\lambda_{\mathrm{y}},\lambda_{\mathrm{z}}\rightarrow 0$ in the limit of infinite pulse duration ($\tau\rightarrow\infty$), as shown in Sec. IV.3.
Finally, we plot numerical results $F\left(U_{\text{ideal}},U\right)$] using the dots in Figs. 4(a)–4(c). The unitary operator $U$ is calculated by numerically solving the Schrödinger equation using Eq. (2). Consequently, the theoretical results are consistent with the numerical analysis, thereby validating the approximations underlying our theoretical analysis.
@@ V.2 Number of control microwave signals NN
We investigate the relationship between the average gate fidelity $F\left(U_{\text{ideal}},U\right)$ and the number of control microwave signals $N$ corresponding to the number of qubits. Figure 7(b) shows $F\left(U_{\text{ideal}},U\right)$ as a function of $N$ for $\tau=\tau_{0}$. An increase in $N$ leads to higher fidelity. The effect of increasing $N$ becomes more pronounced as $k_{0}$ increases. This observed fidelity enhancement is achieved by the deliberate incorporation of suitably chosen off-resonant frequency components into the control pulse. This finding challenges the conventional view that off-resonant microwave components are merely detrimental to qubit control. Therefore, our results align with the emerging understanding of recent research, exemplified by methods such as active leakage cancellation demonstrated in transmon qubits [52].
@@ V.3 Rotation angle $\phi$
VI Conclusion
We have analyzed an FDM-based qubit control scheme that does not require additional circuit elements. We have focused on the pulse length, equivalently considering the spacing of the drive frequencies, to achieve high-fidelity simultaneous gate operations on multiple qubits through a single control line. Our results indicate that orthogonality of the drive-pulse frequencies considerably improves gate fidelity. Notably, even when the frequency spacing was reduced to half of the orthogonal condition, the system retained comparable fidelity levels. These results potentially broaden the design space for spectrum allocation and pulse length of microwave signals.
@@ VI Conclusion
As a final remark, we have neglected leakage to higher energy levels, which can limit gate fidelity in weakly anharmonic systems such as transmons. This simplification allows us to isolate and clarify the essential physics of the OFDM scheme. A full treatment of multilevel dynamics including nonadiabatic transitions is beyond the present scope and will be addressed in future work.

### v2

@@ I Introduction
The developed framework enables the use of closely spaced drive frequencies and short pulse lengths, thereby achieving high throughput in gate operations. Our approach is inherently scalable and well suited for controlling a large number of qubits owing to its simplicity and narrow bandwidth requirements.
@@ II Model
Figure 1: Conceptual illustration of a model comprising NN independent qubits with frequencies $\omega_{\mathrm{q},k}$ driven by microwaves via a shared control line. The controller produces NN microwave tones at distinct frequencies $\omega_{\mathrm{d},k}$ with arbitrary envelopes, which are then combined and routed to a quantum processor via the shared line. The combined signal is applied uniformly to all qubits.
@@
and $\tau$ is the pulse length. In this context, the qubit index set $K$ is given by $K=\left\{l,l+1,\dots,r-1,r\right\}$, where rl+1=Nr-l+1=N. The integers $l\leq 0$ and $r\geq 0$ denote the left and right boundaries of the index set.
@@ III General analysis
by using $\mathrm{sinc}\left(x\right)\equiv\sin\left(x\right)/x$. Here, $\gamma\subset K$ is the set of qubit indices that lack corresponding elements related to $k_{0}$. For example, let $k_{0}=0$. If $2\in K$ while $2\notin K$, then $2\in\gamma$. More formally, a qubit index $k^{\prime}\in K$ belongs to $\gamma$ if and only if its symmetric counterpart with respect to $k_{0}$, namely, $\left(k^{\prime}-k_{0}\right)+k_{0}$, does not belong to $K$ (see Fig. 2).
@@
$\lambda_{\mathrm{z}}$ $\equiv\frac{\alpha^{2}\tau}{4}\left[\sum_{k\in\gamma}\frac{\cos(\Delta_{kk_{0}}\tau)-\frac{3}{2}\mathrm{sinc}(\Delta_{kk_{0}}\tau)}{\Delta_{kk_{0}}}+\sum_{k\in\gamma}\sum_{\begin{subarray}{c}j\in K\\ j\neq k_{0}\end{subarray}}\left\{\frac{\mathrm{sinc}\left(\Delta_{jk_{0}}\tau\right)}{\Delta_{kk_{0}}}+\frac{\mathrm{sinc}\left[\left(\Delta_{kk_{0}}+\Delta_{jk_{0}}\right)\tau\right]}{2\Delta_{jk_{0}}}\right\}\right$.
@@ IV Case studies
Figure 3: Absolute values of the normalized spectra of $s\left(t\right)\sin\left(\omega_{\mathrm{d},k}t\right)$ for $N=5$ and $k\in\left\{-2,-1,0,1,2\right\}$, with pulse length (a) $\tau=\tau_{0}$, (b) $\tau=\tau_{0}/2$, and (c) $\tau=2\tau_{0}$. In (a) and (c), each drive frequency $\omega_{\mathrm{d},k}$ is placed at the zero-crossing point of the spectral components of all other microwaves.
@@ V.2 Number of control microwave signals NN
We investigate the relationship between the average gate fidelity $F\left(U_{\text{ideal}},U\right)$ and the number of control microwave signals $N$ (equal to the number of qubits). Figure 7(b) shows $F\left(U_{\text{ideal}},U\right)$ as a function of $N$ for $\tau=\tau_{0}$. An increase in $N$ leads to higher fidelity. The effect of increasing $N$ becomes more pronounced as $k_{0}$ increases. This observed fidelity enhancement is achieved by the deliberate incorporation of suitably chosen off-resonant frequency components into the control pulse. This finding challenges the conventional view that off-resonant microwave components are merely detrimental to qubit control. Therefore, our results align with the emerging understanding of recent research, exemplified by methods such as active leakage cancellation demonstrated in transmon qubits [52].
@@ V.3 Rotation angle $\phi$
VI Discussion
The pulse length is a crucial factor for enabling dense spacing of drive frequencies while maintaining high gate fidelity. From the results presented in Secs. IV.1IV.2, and V.1, we observe that high fidelity can be attained near $k_{0}=0$ at both $\tau=\tau_{0}$ and $\tau=\tau_{0}/2$. This is because the rotation angles are $\lambda_{\mathrm{x}}\sim\phi/2$ and $\lambda_{\mathrm{y}},\lambda_{\mathrm{z}}\sim 0$ in each case, which is necessary to achieve high fidelity. An equivalent result can be obtained for every integer multiple of $\tau_{0}$ and $\tau_{0}/2$. Consequently, we conclude that $\tau$ should be chosen from the set $\left\{m\tau_{0}/2\right\}$, where $m$ is a positive integer, to obtain higher fidelity around $k_{0}=0$. Note that rectangular pulses with $\tau=m\tau_{0}/2$ can achieve higher gate fidelity than FDM using Gaussian pulses (see Appendix A). In addition, a larger $\tau$ guarantees high fidelity because the rotation angles approach $\lambda_{\mathrm{x}}\rightarrow\phi/2$ and $\lambda_{\mathrm{y}},\lambda_{\mathrm{z}}\rightarrow 0$ in the limit of infinite pulse duration ($\tau\rightarrow\infty$), as shown in Sec. IV.3.
We plot numerical results $F\left(U_{\text{ideal}},U\right)$] using the dots in Figs. 4(a)–4(c). The unitary operator $U$ is calculated by numerically solving the Schrödinger equation using Eq. (2). Consequently, the theoretical results are consistent with the numerical analysis, thereby validating the approximations underlying our theoretical analysis.
Thus far, we have assumed uniform coupling between the drive lines and the qubits. In practice, however, this assumption breaks down: the microwave amplitude experienced by the $k$th qubit can differ from that of others, leading to an incorrect rotation angle and reduced gate fidelity. Although this nonuniform coupling can be partially compensated by adjusting $\alpha_{k}$, the gate fidelity of a qubit tends to be higher when the microwave frequencies are symmetrically allocated around the qubit frequency, as discussed in Sec. V.2. Such compensation therefore introduces asymmetry for the remaining qubits and degrades their fidelities, necessitating a system-level optimization.
[…]
VII Conclusion
We have analyzed an FDM-based qubit control via a single control line. We have focused on the pulse length, equivalently considering the spacing of the drive frequencies, to achieve high-fidelity simultaneous gate operations on multiple qubits. Our results indicate that orthogonality of the drive-pulse frequencies considerably improves gate fidelity. Notably, even when the frequency spacing was reduced to half of the orthogonal condition, the system retained comparable fidelity levels. These results potentially broaden the design space for spectrum allocation and pulse length of microwave signals.
@@ Acknowledgements.
Appendix A Comparison with FDM using Gaussian pulses
(a)(b)(c)
Figure 8: Average gate infidelity $1-F\left(U_{\text{ideal}},U\right)$ as a function of the pulse length $\tau/\tau_{0}$ for Gaussian pulses and rectangular pulses with (a) $k_{0}=0$, (b) $k_{0}=3$, and (c) $k_{0}=6$. Parameters are set as $N=15$, l=r=7-l=r=7, $\omega_{\mathrm{q},k_{0}}/2\pi=5\,\mathrm{GHz}$, $\Delta/2\pi=10\,\mathrm{MHz}$, $\phi=\pi/2$, and $\alpha=\phi/\tau$. The evolution operator $U$ is obtained numerically from the Hamiltonian in Eq. (2).
We compare the infidelity for rectangular pulses under orthogonal and quasi-orthogonal conditions with that for Gaussian pulses. Figures 8(a)–8(c) show the average gate infidelity $1-F\left(U_{\text{ideal}},U\right)$ as a function of the pulse length $\tau/\tau_{0}$ for rectangular pulses and Gaussian pulses with $k_{0}=0$, $3$, and $6$. Here, the evolution operator $U$ is obtained numerically from the Hamiltonian in Eq. (2). For Gaussian pulses, we use the envelope given by [55]
$s\left(t\right)=\phi\frac{\exp\left[-\frac{\left(t-\tau/2\right)^{2}}{2\sigma^{2}}\right]-\exp\left[-\frac{\tau^{2}}{8\sigma^{2}}\right]}{\sqrt{2\pi\sigma^{2}}\mathrm{erf}\left[\frac{\tau}{\sqrt{8}\sigma}\right]-\tau\exp\left[-\frac{\tau^{2}}{8\sigma^{2}}\right]},$ (33)
where $\sigma$ is the standard deviation, and we set $\sigma=\tau/4$. In Figs. 8(a)–8(c), the infidelity for rectangular pulses under orthogonal and quasi-orthogonal conditions ($\tau/\tau_{0}=0.5,1,1.5,2,\dots$) is lower than that for Gaussian pulses with the same or larger values of $\tau$. An exception is observed in Fig. 8(a), where Gaussian pulses yield a lower infidelity in limited regions within $1.5<\tau/\tau_{0}<2.0$. These results indicate that, compared with Gaussian pulses, our approach enables more efficient utilization of the available bandwidth, analogous to OFDM in wireless communication systems.